# VLKEB: A Large Vision-Language Model Knowledge Editing Benchmark

### v2

@@ 1 Introduction
With the rapid advancement and widespread deployment of LLMs, knowledge editing has emerged as an important topic [1, 2]. The knowledge stored within LLMs can suffer from issues such as errors, deficiencies and obsolescence. Knowledge editing aims to efficiently correct and update this information while ensuring minimal impact on unrelated content. Numerous LLM editing methods have been proposed [3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and benchmarks have been established to assess these approaches. Metrics such as Reliability, Generality, Locality and Portability are commonly used [2] in these benchmarks. Reliability defines a reliable edit where the post-edit model produces the target answer for a given case. Generality requires the post-edit model to understand equivalent neighbors, such as rephrased sentences. Locality emphasizes localized editing, ensuring that the output of unrelated knowledge is retained. Portability evaluates whether models effectively apply edited knowledge to relevant content. These methods and benchmarks have greatly advanced the research in LLM knowledge editing.
@@
In our study, we source data from the multi-modal knowledge graph MMKG  [15], which includes images linked to knowledge entities. The presence of multiple images for each entity in MMKG makes it practicable to selectively choose image pairs for assessing the Generality. During the selection process, we choose clear and representative image pairs, ensuring that each pair contains a same entity but presented in varied perspectives or appearances. For the Reliability test, corresponding entities of chosen images are employed. Subsequently, we filter similar images from the remaining entities (none of their images were chosen previously) to construct the image Locality test. We utilize ChatGPT APIs to generate questions and answers for these entities and images during data construction.
@@ 4.2 Single Editing Results and Findings
(Rel.: Reliability; T/I-Gen.: Text/Image Generality; T/I-Loc.: Text/Image Locality; Port.: Portability)
@@ B.1 Metrics
$\mathcal{M}_{\text{generality}}^{\text{text}}=\underset{\begin{subarray}{c}\left(i_{e},x_{e},y_{e},y_{e}^{\prime}\right)\sim\mathcal{D}_{\text{edit }}\\ x_{r}\sim\mathcal{N}\left(x_{e}\right)\end{subarray}}{\mathbb{E}}\mathbbm{1}\left\{f\left(i_{e},x_{r};\theta^{\prime}\right)=y_{e}\right\}$
and
@@
$\mathcal{M}_{\text{locality}}^{\text{text}}=\underset{\left(x_{l},y_{l}\right)\sim\mathcal{D}_{\text{loc}}^{\text{text}}}{\mathbb{E}}\mathbbm{1}\left\{f\left(x_{l};\theta^{\prime}\right)=f(x_{1};\theta)\right\}$
and
@@ B.4 Experiment Resources and Parameters
In this study, we utilize an internal cluster equipped with the following resources: AMD EPYC 7763 CPUs, NVIDIA A100 80GB GPUs, and 512GB of RAM. The operating system is Ubuntu 20.04.6, and we employ PyTorch in our experiments. The parameters are in the following and listed in config files in code repository.
@@
We use the sentence-transformers model all-MiniLM-L6-v2 to embed texts and retrieve similar edits in training set. The number of demonstrations is set to 32 for all models.
@@ C.2 Prompt Template for QA Generation
[…]
@@
[…]
@@ C.3 Examples of Each Test
Q: Who is the actor featured in this image? Target: Andre Braugher
@@
Q: What city is the birth place of the person in the picture? A: Chicago
@@
Q: Where is the alma mater of the person associated with the person in the picture located? A: New York City
@@
Q: What is the time zone of the city where the alma mater of the person connected to the individual in the picture is located? A: Eastern Time Zone
@@
Q: What is the time zone of the largest city in the country where the residence of the person associated with the entity shown in the picture is located? A: Eastern Time Zone
@@
Edit Q: In which city was Andre Braugher born? Target: Chicago
@@
Test Q: What city is the birth place of the person in the picture? A: Chicago
@@ Appendix D Original Results of Multi-hop Portability and Sequential Editing
In Sec.4.3, we present the results of multi-hop Portability in figures. Here we provide the original results in Tab.12.
In Sec.4.4, we present the results of sequential editing in figures. Here we provide the original results in Tab.LABEL:tab:seq_edit.

### v3

@@ 1 Introduction
With the rapid advancement and widespread deployment of LLMs, knowledge editing has emerged as an important topic [1, 2, 3, 4]. The knowledge stored within LLMs can suffer from issues such as errors, deficiencies and obsolescence. Knowledge editing aims to efficiently correct and update this information while ensuring minimal impact on unrelated content. Numerous LLM editing methods have been proposed [5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and benchmarks have been established to assess these approaches. Metrics such as Reliability, Generality, Locality and Portability are commonly used [2] in these benchmarks. Reliability defines a reliable edit where the post-edit model produces the target answer for a given case. Generality requires the post-edit model to understand equivalent neighbors, such as rephrased sentences. Locality emphasizes localized editing, ensuring that the output of unrelated knowledge is retained. Portability evaluates whether models effectively apply edited knowledge to relevant content. These methods and benchmarks have greatly advanced the research in LLM knowledge editing.
@@
In our study, we source data from the multi-modal knowledge graph MMKG  [17], which includes images linked to knowledge entities. The presence of multiple images for each entity in MMKG makes it practicable to selectively choose image pairs for assessing the Generality. During the selection process, we choose clear and representative image pairs, ensuring that each pair contains a same entity but presented in varied perspectives or appearances. For the Reliability test, corresponding entities of chosen images are employed. Subsequently, we filter similar images from the remaining entities (none of their images were chosen previously) to construct the image Locality test. We utilize GPT to generate questions and answers during data construction.
@@ 3.1 Metrics: Reliability, Generality, Locality and Portability
where $x_{p}$ and $y_{p}$ are text inputs and outputs that belong to $\mathcal{P}(i_{e},x_{e},y_{e},y_{e}^{\prime})$, which denotes the portability scope given the input $i_{e},x_{e},y_{e}$, and target output $y_{e}^{\prime}$.
@@ 4.2 Single Editing Results and Findings
Rel.: Reliability; T/I-Gen.: Text/Image Generality; T/I-Loc.: Text/Image Locality; Port.: Portability
@@ 5 Conclusion, Limitation and Future Direction
Acknowledgments and Disclosure of Funding
This work is jointly sponsored by National Natural Science Foundation of China (62236010, 62141608, 62206291).
@@ B.1 Metrics
$\mathcal{M}_{\text{generality}}^{\text{text}}=\underset{\begin{subarray}{c}\left(i_{e},x_{e},y_{e},y_{e}^{\prime}\right)\sim\mathcal{D}_{\text{edit }}\\ x_{r}\sim\mathcal{N}\left(x_{e}\right)\end{subarray}}{\mathbb{E}}\mathbbm{1}\left\{f\left(i_{e},x_{r};\theta^{\prime}\right)=y_{e}\right\},$
@@
$\mathcal{M}_{\text{locality}}^{\text{text}}=\underset{\left(x_{l},y_{l}\right)\sim\mathcal{D}_{\text{loc}}^{\text{text}}}{\mathbb{E}}\mathbbm{1}\left\{f\left(x_{l};\theta^{\prime}\right)=f(x_{1};\theta)\right\},$
@@ B.4 Experiment Resources and Parameters
In this study, we utilize an internal cluster equipped with NVIDIA A100 80GB GPUs, and we employ PyTorch in our experiments. The parameters are in the following and in config files in code repository.
@@
Use the sentence-transformers model all-MiniLM-L6-v2 to embed texts and retrieve similar edits in training set. The number of demonstrations is set to 32 for all models.
@@ C.2 Prompt Template for QA Generation
Q1: What university is featured in the image?
Q2: Which educational institution is depicted in the picture?
A: University of New Mexico.
@@
Q: What is the country associated with the birthplace of the spouse of the person shown in the picture?
A: English language
@@ C.3 Examples of Each Test
Q: Who is the actor featured in this image?
Target: Andre Braugher
@@
Q: What city is the birth place of the person in the picture?
A: Chicago
@@
Q: Where is the alma mater of the person associated with the person in the picture located?
A: New York City
@@
Q: What is the time zone of the city where the alma mater of the person connected to the individual in the picture is located?
A: Eastern Time Zone
@@
Q: What is the time zone of the largest city in the country where the residence of the person associated with the entity shown in the picture is located?
A: Eastern Time Zone
@@
Edit Q: In which city was Andre Braugher born?
Target: Chicago
@@
Test Q: What city is the birth place of the person in the picture?
A: Chicago
@@ Appendix D Original Results of Multi-hop Portability and Sequential Editing
In Sec.4.3, we present the results of multi-hop Portability in figures. Here we provide the original results in Tab.12. In Sec.4.4, we present the results of sequential editing in figures. Here we provide the original results in Tab.LABEL:tab:seq_edit.